OpenStudy (thatgirl21):

Please help explain how to solve! The income tax, T, owed in a certain country given that the income is I is given by the formula: T(i)=0.18i - 5,000 a.) what is the income tax owed on $60,000 b.) how much income is needed to generate $16,600 worth of taxes?

OpenStudy (maddyhall15):

a) $5800 Plug in $60000 into the formula. T(i)=0.18(60000)-5000 T(i)=5800 b) $95000 replace T(i) with $16600 solve for i. 16600=0.18i-5000 +5000 +5000 17100/0.18=0.18i/0.18 95000=i

OpenStudy (thatgirl21):

Thank you so much!!